# Gaussian Splatting with Localized Points Management

### v2

@@ 4 Experiment
We conducted an extensive evaluation using both static and dynamic scenes derived from publicly datasets. For static scenes, our approach was applied to a total of 11 scenes as specified in the 3DGS framework [13], which includes nine scenes from Mip-NeRF360 [3], two from Tanks&Temples [14], and two from DeepBlending [12]. In the context of dynamic scenes, our approach was tested across six scenes from the Neural 3D Video Dataset [16].
To evaluate novel view synthesis performance, we follow standard protocols by selecting one out of every eight images as test images, with the remaining used for training in static scenes. For each dynamic scene within the Neural 3D Video Dataset, one view was designated for testing while the others were allocated for training purposes. Evaluation metrics included the peak signal-to-noise ratio (PSNR), structural similarity index measure (SSIM), and the learned perceptual image patch similarity (LPIPS), which are broadly recognized standards in the field.
Baseline and Implementation
Vanilla 3D Gaussian Splatting(3DGS) [13] and its variants SpacetimeGS(STGS)  [18] are selected as our main baseline for its established SOTA performance in novel view synthesis. For the Static 3D benchmark, We also record the results of Mip-NeRF360 [3], iNGP  [25] and Plenoxels  [11] as in  [13]. For the Dynamic 4D benchmark, we also performed system comparison, such as DyNeRF [17], StreamRF [15], K-planes [10] and so on. In alignment with the approach described in 3DGS an STGS, our models are trained for 30k iterations across all scenes, following the same training schedule and hyperparameters. In addition to the original Gaussian densification strategies used in 3DGS and SpaceTime Gaussian, we also perform localized points management, including addition, reset, and pruning. We maintain the same thresholds for splitting and cloning points as in the original 3DGS and SpaceTime Gaussian. For point matching, we perform offline extraction to save computational cost. All experiments were conducted on an RTX 3090 GPU with 24GB of memory.
Figure 3: Qualitative evaluation of our LPM across diverse static datasets [4, 12, 14]. Our LPM improves 3DGS [13] on these challenging scenarios, e.g. (a) Light effect, (b) Completeness in the distance, (c) Intricate details and (d) Transparent. Patches that highlight the visual differences are emphasized with red insets for clearer visibility.
@@ 4.1 Main Results
Method Mip-NeRF 360 Tanks&Temples Deep Blending
PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS
Plenoxels  [11] 23.08 0.625 0.463 21.08 0.719 0.379 23.06 0.795 0.510
INGP-Big  [25] 25.59 0.699 0.331 21.92 0.745 0.305 24.96 0.817 0.390
Mip-NeRF 360  [3] 27.69 0.792 0.237 22.22 0.759 0.257 29.40 0.901 0.245
3DGS  [13] 27.21 0.815 0.214 23.14 0.841 0.183 29.41 0.903 0.243
3DGS* 27.47 0.816 0.216 23.67 0.849 0.177 29.55 0.904 0.245
3DGS + LPM 27.59 0.820 0.216 23.83 0.850 0.181 29.76 0.908 0.241
The quantitative results (PSNR, SSIM, and LPIPS) on the Mip-NeRF 360 and Tanks & Temples datasets are presented in Tables 1. We retrained the 3DGS model (referred to as 3DGS*) as it yields better performance compared to the original 3DGS. Our approach achieves results comparable to the state-of-the-art on the Mip-NeRF360 dataset and further enhances 3DGS using our point management technique. Additionally,  LPM improve 3DGS to set new state-of-the-art results on the Tanks & Temples and DeepBlending datasets, effectively capturing more challenging environments (e.g., light effects, transparency). These results quantitatively validate the effectiveness of our method in improving the quality of reconstruction.
In Figures 3, we present a comparison between 3DGS [13] and the 3DGS+LPM. A variety of improvements can be observed, particularly in challenging cases such as light effects, completeness at a distance, intricate details, and transparency. Our LPM significantly reduces artifacts in specific regions on top of 3DGS, particularly in the tree and flowers in the second and third rows. These regions require more points for accurate population, leading to a more precise and detailed reconstruction. Additionally, the tablecloth and window regions in the first and last rows are affected by ill-conditioned points. Our geometry calibration with LPM provides an opportunity to correct these potentially ill-conditioned points, enhancing the overall reconstruction accuracy.
@@
Method PSNR DSSIM1 DSSIM2 LPIPS FPS
LLFF [23] 1 23.24 - 0.076 0.235 -
DyNeRF [17] 1 29.58 0.020 0.083 0.063 0.015
Dynamic-4DGS [36] 1 - - - - 30
4DGS  [40] 1 29.38 - - - 114
STGS  [18] 1 29.58 0.038 0.022 0.063 103
STGS* 1 29.48 0.038 0.023 0.066 110
STGS 1 + LPM 29.84 0.036 0.022 0.062 105
StreamRF [15] 28.26 - - 0.039 10.9
NeRFPlayer [29] 30.69 0.034 - 0.111 0.05
HyperReal [2] 31.10 0.036 - 0.096 2
K-planes [10] 31.63 0.018 - 0.31 3
MixVoxels-X  [34] 31.73 0.015 - 0.064 4.6
Dynamic-4DGS  [36] 31.15 - 0.016 0.049 30
4DGS  [40] 32.01 - - 0.055 114
STGS  [18] 32.05 0.026 0.014 0.044 140
STGS* 31.99 0.026 0.015 0.045 145
STGS+ LPM 32.40 0.025 0.014 0.045 140
Table  2 presents a quantitative evaluation on the Neural 3D Video Dataset. Following established practices, training and evaluation are conducted at half resolution, with the first camera held out for evaluation  [17]. Integrating our LPM into SpaceTimeGS yields the best performance across all comparisons. Notably, our method demonstrates significant improvements in the challenging Flame Salmon scene compared to STGS [18]. Our approach not only surpasses previous methods in rendering quality but also maintains comparable rendering speed.
@@ 4.2 Ablation study
Effectiveness and efficiency of the LPM
We hypothesize that the Adaptive Density Control (ADC) tends to overlook under-optimized points due to its simplistic approach of thresholding the average gradient. The straight way to identify the all points is lowering threshold to densification process. Although this solution can reduce blurring in specific regions, such as the toy (red box) illustrated in Figure 5, it still has limitations. As shown in Table 3, lowering the threshold for 3DGS significantly increases the number of Gaussian points and decreases rendering speed. Additionally, the PSNR of the quantitative results decreases due to the introduction of unnecessary points in already dense areas. In contrast, LPM effectively generates points in areas indicated by the error map, leading to more accurate and detailed reconstructions while maintaining real-time rendering speed. As demonstrated by the qualitative comparison in Figure 5, 3DGS with  LPM achieves superior qualitative results.
Table 3: Efficiency analysis. Rendering speed of both methods are measured on our machine.
Scene Method PSNR LPIPS Gaussians Training time
3DGS*  [13] 30.03 0.244 232k 22min
PlayRoom 3DG* (lower threshold) 29.69 0.240 523k 36min
3DGS + LPM 30.22 0.241 186k 23min
3DGS* [13] 25.42 0.146 257k 19 min
Truck 3DGS* (lower threshold) 25.45 0.127 635k 35min
3DGS + LPM 25.61 0.154 265k 21min
Individual points manipulation
Table 4: Performance comparison for different configurations
@@
5 Conclusions and limitations
Appendix A Appendix
A.1 Additional Results
@@
Plenoxels [11] 21.912 20.097 23.4947 20.661 22.487 27.594
INGP-Big [25] 22.171 20.652 25.069 23.466 22.373 29.690
Mip-NeRF 360 [3] 24.37 21.73 26.98 26.40 22.87 31.63
3DGS  [13] 25.246 21.520 27.410 26.550 22.490 30.632
@@
Plenoxels [11] 23.624 23.420 24.669 23.142 22.980 23.221 18.927
INGP-Big  [25] 26.691 29.479 30.685 28.257 21.665 23.383 20.456
Mip-NeRF 360  [3] 29.55 32.23 33.46 29.140 29.657 24.912 19.523
3DGS  [13] 28.700 30.317 31.980 28.766 30.044 25.187 21.097
@@
Plenoxels  [11] 0.506 0.521 0.3864 0.503 0.540 0.4186
INGP-Big  [25] 0.446 0.441 0.257 0.421 0.450 0.261
Mip-NeRF 360 [3] 0.301 0.344 0.170 0.261 0.339 0.211
3DGS  [13] 0.205 0.336 0.103 0.210 0.317 0.220
@@
Plenoxels  [11] 0.441 0.447 0.398 0.521 0.499 0.335 0.422
INGP-Big  [25] 0.306 0.195 0.205 0.352 0.428 0.249 0.360
Mip-NeRF 360  [3] 0.204 0.127 0.176 0.237 0.252 0.159 0.354
3DGS  [13] 0.204 0.129 0.205 0.244 0.241 0.148 0.218
@@
Plenoxels  [11] 0.496 0.431 0.6063 0.523 0.509 0.8417
INGP-Big [25] 0.512 0.486 0.701 0.594 0.542 0.871
Mip-NeRF 360  [3] 0.685 0.583 0.813 0.744 0.632 0.913
3DGS  [13] 0.771 0.605 0.868 0.775 0.638 0.914
@@
Plenoxels  [11] 0.759 0.648 0.814 0.787 0.802 0.774 0.663
INGP-Big  [25] 0.817 0.858 0.906 0.854 0.779 0.800 0.689
Mip-NeRF 360  [3] 0.894 0.920 0.941 0.901 0.900 0.857 0.660
3DGS  [13] 0.905 0.922 0.938 0.899 0.906 0.879 0.802
@@
In Table  12, we provide the PSNR on different scenes. The quanlitative results clearly show that LPM improve STGS  [18] to faithfully capture the subtle static and dynamic information.
@@
K-Planes-explicit  [10] 28.74 32.19 31.93 28.71 31.80 31.89
K-Planes-hybrid  [10] 29.99 32.60 31.82 30.44 32.38 32.52
MixVoxels  [34] 29.36 31.61 31.30 29.92 31.21 31.43
NeRFPlayer  [29] 31.53 30.56 29.35 31.65 31.93 29.12
HyperReel  [2] 28.37 32.30 32.92 28.26 32.20 32.57
Dynamic-4D  [36] 27.34 32.46 32.90 29.20 32.51 32.49
4DGS  [40] 28.33 32.93 33.85 29.38 34.03 33.51
STGS  [18] 28.61 33.18 33.52 29.48 33.64 33.89
@@ Appendix B More visualizations
Figure 6: More qualitative comparisons on static 3D and dynamic 4D dataset.

### v3

@@ 4 Experiment
We conducted an extensive evaluation using both static and dynamic scenes derived from publicly datasets. For static scenes, our approach was applied to a total of 13 scenes as specified in the 3DGS framework [14], which includes nine scenes from Mip-NeRF360 [3], two from Tanks&Temples [15], and two from DeepBlending [12]. In the context of dynamic scenes, our approach was tested across six scenes from the Neural 3D Video Dataset [17].
To evaluate novel view synthesis performance, we followed standard protocols by selecting one out of every eight images as test images, with the remaining used for training in static scenes. For each dynamic scene within the Neural 3D Video Dataset, one view was designated for testing while the others were allocated for training purposes. Evaluation metrics included the peak signal-to-noise ratio (PSNR), structural similarity index measure (SSIM), and the learned perceptual image patch similarity (LPIPS), which are broadly recognized standards in the field.
Baselines and implementation
Vanilla 3D Gaussian Splatting (3DGS) [14], 2D Gaussian Splatting (2DGS)  [13], Mip Gaussian Splatting (MipGS) [40], PiexlGS  [41] and SpacetimeGS (STGS) [18] were selected as our main baselines for their established art performance in novel view synthesis. For the static 3D benchmark, we also recorded the results of Mip-NeRF360 [3], iNGP [24] and Plenoxels  [10] as in [14]. For the Dynamic 4D benchmark, we performed system comparison, such as DyNeRF [16], K-planes [11] and so on. In alignment with the approach described in 3DGS an STGS, our models were trained for 30k iterations across all scenes, following the same training schedule and hyperparameters. In addition to the original Gaussian densification strategies used in 3DGS and SpaceTime Gaussian, we also performed localized points management, including addition, reset, and pruning. We maintained the same thresholds for splitting and cloning points as in the original 3DGS and SpaceTime Gaussian. For point matching, we performed offline extraction to save computational cost. All experiments were conducted on an RTX 3090 GPU with 24GB of memory.
@@ 4.1 Main Results
Method PSNR DSSIM LPIPS FPS
LLFF 1 23.24 0.076 0.235 -
DyNeRF 1 29.58 0.083 0.063 0.015
Dynamic-4DGS 1 - - - 30
4DGS 1 29.38 - - 114
STGS 1 29.58 0.022 0.063 103
STGS* 1 29.48 0.023 0.066 110
STGS* 1 + LPM 29.84 0.022 0.062 105
StreamRF 28.26 - 0.039 10.9
NeRFPlayer 30.69 - 0.111 0.05
HyperReal 31.10 - 0.096 2
K-planes 31.63 - 0.31 3
MixVoxels-X 31.73 - 0.064 4.6
Dynamic-4DGS 31.15 0.016 0.049 30
4DGS 32.01 - 0.055 114
STGS 32.05 0.014 0.044 140
STGS* 31.99 0.015 0.045 145
STGS*+ LPM 32.40 0.014 0.045 140
The quantitative results (PSNR, SSIM, and LPIPS) on the Mip-NeRF 360 and Tanks & Temples datasets are presented in Tables 1. We retrained the 3DGS model (referred to as 3DGS*) as it yields better performance compared to the vanilla 3DGS and its variants. Our approach achieves results comparable to the state-of-the-art on the Mip-NeRF360 dataset and further enhances all 3DGS based method using our point management technique. Additionally,  LPM improve vanilla 3DGS and PiexlGS to set new state-of-the-art results on the Mip-NeRF 360, Tanks & Temples datasets and Dep Blending, effectively capturing more challenging environments (e.g., light effects, transparency). These results quantitatively validate the effectiveness of our method in improving the quality of reconstruction.
In Figures 3, we present a comparison between 3DGS [14] and 3DGS* + LPM, focusing on both appearance and depth. A variety of improvements can be observed, particularly in challenging cases such as light effects, completeness at a distance. Our LPM significantly reduces artifacts in specific regions on top of 3DGS, particularly in the tree at the second. These regions require more points for accurate population, leading to a more precise and detailed reconstruction. Additionally, the tablecloth in the first row is affected by ill-conditioned points. Furthermore, we provide depth and mesh comparisons in the third and final rows. All these observations demonstrate that our geometry calibration with LPM offers an opportunity to correct these potentially ill-conditioned points, thereby enhancing the overall reconstruction accuracy.
@@
Table  2 presents a quantitative evaluation on the Neural 3D Video Dataset. Following established practices, training and evaluation are conducted at half resolution, with the first camera held out for evaluation  [16]. Integrating our LPM into SpaceTimeGS yields the best performance across all comparisons. Notably, our method demonstrates significant improvements in the challenging Flame Salmon scene compared to SpaceTimeGS  [18]. Our approach not only surpasses previous methods in rendering quality but also maintains comparable rendering speed.
@@ 4.2 Ablation study
Table 3: Performance comparison for different configurations
Effectiveness and cost of LPM
Scene Method PSNR LPIPS Gaussians Training time
3DGS* 30.03 0.244 232k 22min
3DG* (lower threshold) 29.69 0.240 523k 36min
PlayRoom GaussianPro Out of Memory
PiexlGS 30.09 0.241 186k 35min
3DGS + LPM 30.22 0.241 186k 23min
3DGS* 25.42 0.146 257k 19 min
3DGS* (lower threshold) 25.45 0.127 635k 35min
Truck GaussianPro 25.40 0.164 312k 36min
PiexlGS 25.51 0.121 518k 37min
3DGS + LPM 25.61 0.154 265k 21min
Table 4: Cost-effectiveness analysis. Rendering speed of both methods are measured on our machine. Note: For 3DGS+LPM, training time includes the feature matching process.
We hypothesize that the Adaptive Density Control (ADC) tends to overlook under-optimized points due to its simplistic approach of thresholding the average gradient. The straight way to identify the all points is lowering threshold to densification process. Although this solution can reduce blurring in specific regions, such as the toy (red box) illustrated in Figure 4, it still has limitations. As shown in Table 4, lowering the threshold for 3DGS significantly increases the number of Gaussian points and decreases rendering speed. Additionally, the PSNR of the quantitative results decreases due to the introduction of unnecessary points in already dense areas. In contrast, LPM effectively generates points in areas indicated by the error map, leading to more accurate and detailed reconstructions while maintaining real-time rendering speed. As demonstrated by the qualitative comparison in Figure 4, 3DGS with  LPM achieves superior qualitative results. We further compare our method with other recent methods that also focus on adaptive density control (ADC). While PixelGS and GaussianPro achieve improvements in rendering quality, their training times increase substantially as they only consider point addition and extra gradient propagation. In contrast, LPM achieves a noticeable improvement with only a slight increase in training time due to (1) point matching [19] is much faster (2) considering model expansion to dynamically prune the points by their addition number and (3) selecting points in error-contributing zones 3D zone using the parallel matrix operations.
Individual points manipulation
5 Conclusion
Supplementary Material
6 Additional Results
@@
Plenoxels 21.912 20.097 23.4947 20.661 22.487 27.594
INGP-Big 22.171 20.652 25.069 23.466 22.373 29.690
Mip-NeRF 360 24.37 21.73 26.98 26.40 22.87 31.63
3DGS 25.246 21.520 27.410 26.550 22.490 30.632
@@
Plenoxels 23.624 23.420 24.669 23.142 22.980 23.221 18.927
INGP-Big 26.691 29.479 30.685 28.257 21.665 23.383 20.456
Mip-NeRF 360 29.55 32.23 33.46 29.140 29.657 24.912 19.523
3DGS 28.700 30.317 31.980 28.766 30.044 25.187 21.097
@@
Plenoxels 0.506 0.521 0.3864 0.503 0.540 0.4186
INGP-Big 0.446 0.441 0.257 0.421 0.450 0.261
Mip-NeRF 360 0.301 0.344 0.170 0.261 0.339 0.211
3DGS 0.205 0.336 0.103 0.210 0.317 0.220
@@
Plenoxels 0.441 0.447 0.398 0.521 0.499 0.335 0.422
INGP-Big 0.306 0.195 0.205 0.352 0.428 0.249 0.360
Mip-NeRF 360 0.204 0.127 0.176 0.237 0.252 0.159 0.354
3DGS 0.204 0.129 0.205 0.244 0.241 0.148 0.218
@@
Plenoxels 0.496 0.431 0.6063 0.523 0.509 0.8417
INGP-Big 0.512 0.486 0.701 0.594 0.542 0.871
Mip-NeRF 360 0.685 0.583 0.813 0.744 0.632 0.913
3DGS 0.771 0.605 0.868 0.775 0.638 0.914
@@
Plenoxels 0.759 0.648 0.814 0.787 0.802 0.774 0.663
INGP-Big 0.817 0.858 0.906 0.854 0.779 0.800 0.689
Mip-NeRF 360 0.894 0.920 0.941 0.901 0.900 0.857 0.660
3DGS 0.905 0.922 0.938 0.899 0.906 0.879 0.802
@@
Method Indoor Outdoor
PSNR SSIM LPIPS PSNR SSIM LPIPS
2DGS* 24.210 0.705 0.282 30.105 0.911 0.211
2DGS + LPM 24.427 0.716 0.264 30.432 0.919 0.193
@@
In Table  13, we provide the PSNR on different scenes. The quanlitative results clearly show that LPM improve STGS  [18] to faithfully capture the subtle static and dynamic information.
@@
K-Planes-explicit 28.74 32.19 31.93 28.71 31.80 31.89
K-Planes-hybrid 29.99 32.60 31.82 30.44 32.38 32.52
MixVoxels 29.36 31.61 31.30 29.92 31.21 31.43
NeRFPlayer 31.53 30.56 29.35 31.65 31.93 29.12
HyperReel 28.37 32.30 32.92 28.26 32.20 32.57
Dynamic-4D 27.34 32.46 32.90 29.20 32.51 32.49
4DGS 28.33 32.93 33.85 29.38 34.03 33.51
STGS 28.61 33.18 33.52 29.48 33.64 33.89
@@ 7 More visualizations
Figure 6: Additional qualitative comparisons on static 3D and dynamic 4D datasets.